Let ~u =< −2, 1, −1, ~v =< −3, 2, −1 > and w~ =< 1, 3, 5 >. Compute:

(6.1) ~u × w~ ,

(6.2) ~u × (w~ × ~v) and (~u × w~ ) × ~v.


1
Expert's answer
2021-06-27T18:38:50-0400

(6.1)


"\vec u\times \vec w=\begin{vmatrix}
 \vec i & \vec j & \vec k \\
 -2 & 1 & -1 \\
 1 & 3 & 5 \\
\end{vmatrix}"

"=\vec i\begin{vmatrix}
 1 & -1 \\
 3 & 5
\end{vmatrix}-\vec j\begin{vmatrix}
 -2 & -1 \\
 1 & 5
\end{vmatrix}+\vec k\begin{vmatrix}
 -2 & 1 \\
 1 & 3
\end{vmatrix}"

"=8\vec i+9\vec j-7\vec k"

"\vec u\times \vec w=\langle8, 9, -7\rangle"

(6.2)

"\vec w\times \vec v=\begin{vmatrix}
 \vec i & \vec j & \vec k \\
 1 & 3 & 5 \\
 -3 & 2 & -1 \\
\end{vmatrix}"

"=\vec i\begin{vmatrix}
 3 & 5 \\
 2 & -1
\end{vmatrix}-\vec j\begin{vmatrix}
 1 & 5 \\
 -3 & -1
\end{vmatrix}+\vec k\begin{vmatrix}
 1 & 3 \\
 -3 & 2
\end{vmatrix}"

"=-13\vec i-14\vec j+11\vec k"

"\vec u\times (\vec w\times \vec v)=\begin{vmatrix}
 \vec i & \vec j & \vec k \\
 -2 & 1& -1 \\
 -13 & -14 & 11 \\
\end{vmatrix}"

"=\vec i\begin{vmatrix}
 1 & -1 \\
 -14 & 11
\end{vmatrix}-\vec j\begin{vmatrix}
 -2 & -1 \\
 -13 & 11
\end{vmatrix}+\vec k\begin{vmatrix}
 -2 & 1 \\
 -13 & -14
\end{vmatrix}"

"=-3\vec i+35\vec j+41\vec k"

"\vec u\times (\vec w\times \vec v)=\langle -3, 35, 41 \rangle"


"(\vec u\times \vec w)\times \vec v=\begin{vmatrix}
 \vec i & \vec j & \vec k \\
 8 & 9 & -7 \\
 -3 & 2 & -1 \\
\end{vmatrix}"

"=\vec i\begin{vmatrix}
 9 & -7 \\
 2 & -1
\end{vmatrix}-\vec j\begin{vmatrix}
 8 & -7 \\
 -3 & -1
\end{vmatrix}+\vec k\begin{vmatrix}
 8 & 9 \\
 -3 & 2
\end{vmatrix}"

"=5\vec i+29\vec j+43\vec k"

"(\vec u\times \vec w)\times \vec v=\langle 5, 29, 43 \rangle"